# Train for the Worst, Plan for the Best: Understanding Token Ordering in Masked Diffusions

### v2

@@ D.1.1 Experiment on L&O-NAE-SAT distribution
We consider five instances of L&O-NAE-SAT: $(N,P)=(25,275),(30,270),(40,260),(50,250),(100,200)$. For each distribution, we train a 19M MDM and measure the accuracy difference between vanilla inference and adaptive inference using Top-K probability margin.
@@ D.1.2 Experiment on text data
Top-$K$ probability margin sampler with temperature.
To modify our inference for text data modeling, which does not have a determined answer, we found that adding a certain level of temperature to the oracle is useful. This is because Top-$K$ probability margin or Top-$K$ often leads to greedy sampling, which harms the diversity (entropy) of the generated samples. Therefore, we consider a variant of the oracle as follows, incorporating a noise term $\epsilon$:
@@ D.2 Experimental details on Sudoku and Zebra puzzles
For the Sudoku dataset, we use $6$M GPT-2 model and for the Zebra dataset, we use $19$M model but instead of causal attention, we use complete bidirectional attention. We set the learning rate to 0.001 with batch size 128 to train the model for 300 epochs. For the inference, we use 50 reverse sampling steps using the appropriate strategy. Additionally, we add Gumbel noise with a coefficient of $0.5$ to the MDM inference oracle $\mathcal{F}$.
@@ Appendix E Omitted proofs
We first re-state the Proposition 3.1 from (Zheng et al., 2024). To clarify, (Zheng et al., 2024) generally considers the case beyond the time-embedding denoising network $p_{\theta}$.
@@
For clean data $x_{0}$, let $\tilde{q}(x(n)\mid x_{0})$ be the discrete forward process that randomly and uniformly masks $n$ tokens of $x_{0}$. Suppose $\alpha_{0}=0$ and $\alpha_{1}=1$. Then the MDM training loss (1) can be reformulated as
$\mathcal{L}_{\theta}=-\sum_{n=1}^{L}\mathbb{E}_{x(n)\sim\tilde{q}(\cdot\mid x_{0})}\left[\frac{1}{n}\sum_{\ell:x(n)^{\ell}=m}\mathbf{e}_{x_{0}^{\ell}}\log p_{\theta}(x^{\ell}\mid x(n))\right].$ (6)
To obtain an alternative formulation of (6), we expand the expectation $x(n)\sim\tilde{q}(\cdot\mid x_{0})$. Since there are total $L$ positions of $x_{0}$, we have the probability assigned for each $x(n)$ equals $1/\binom{L}{n}$. Therefore,
$\mathcal{L}_{\theta}=$ $-\sum_{n=1}^{L}\mathbb{E}_{x(n)\sim\tilde{q}(\cdot\mid x_{0})}\left[\frac{1}{n}\sum_{\ell:x(n)^{\ell}=m}\mathbf{e}_{x_{0}^{\ell}}\log p_{\theta}(x^{\ell}\mid x(n))\right]$
$=-\sum_{M\in[L],i\in M}\frac{1}{\binom{L}{|M|}}\times\frac{1}{|M|}\mathbf{e}_{x_{0}^{\ell}}\log p_{\theta}(x^{\ell}\mid x[M])$
$=-\sum_{M\in[L],i\in M}\frac{1}{\binom{L}{|M|}}\times\frac{1}{|M|}\log p_{\theta}(x_{0}^{\ell}\mid x[M])$
$=-\sum_{M\in[L],i\in M}\frac{1}{L\binom{L-1}{|M|-1}}\log p_{\theta}(x_{0}^{\ell}\mid x[M]).$
Reformulating the MDM loss with $\pi$-learner s..
In this paragraph, we provide the proof of
$-\frac{1}{L}\sum_{M\subseteq[L],i\in M}\frac{1}{\binom{L-1}{|M|-1}}\displaystyle\mathop{\mathbb{E}}_{x_{0}\sim p_{\textrm{d}ata}}[\log p_{\theta}(x^{i}_{0}|x_{0}[M])]=-\mathbb{E}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L}),x_{0}\sim p_{\rm{data}}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right].$
Alternatively, we will demonstrate that
$-\frac{1}{L}\sum_{M\subseteq[L],i\in M}\frac{1}{\binom{L-1}{|M|-1}}\displaystyle\log p_{\theta}(x^{i}_{0}|x_{0}[M])=-\mathbb{E}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L})}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$
holds for every $x_{0}$. Note that
$\mathbb{E}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L})}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$
$=$ $\frac{1}{L!}\sum_{\pi\in\mathbb{S}_{L}}\sum_{j=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(j)}\Big{|}x_{0}[\pi\{j,\ldots,L-1\}]\right).$
Next, by regarding $\pi\{j,\dots,L-1\}=\{\pi(j),\dots,\pi(L-1)\}=M\subseteq[L]$ and $\pi(j)=i$ in the equation  (1), we count the number of $\pi\in S_{L}$ that induces a specific term $\log p_{\theta}(x_{0}^{i}|x_{0}[M])$. For a given $M\in[L]$ and $i\in M$, $\pi$ must satisfy
@@
The number of $\pi$ that satisfies above is $(L-|M|)!\times(|M|-1)!$. Finally, the following calculation concludes the proof.
$\mathbb{E}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L})}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$
$=$ $\frac{1}{L!}\sum_{\pi\in S_{L}}\sum_{j=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(j)}\Big{|}x_{0}[\pi\{j,\ldots,L-1\}]\right)$
$=$ $\frac{1}{L!}\sum_{|M|\in[L],i\in M}\big{[}\log p_{\theta}(x_{0}^{i}|x_{0}[M])\times(L-1-|M|)!\times(|M|-1)!\big{]}$
$=$ $\frac{1}{L}\sum_{|M|\in[L],i\in M}\frac{1}{\binom{L-1}{|M|-1}}\times\log p_{\theta}(x_{0}^{i}|x_{0}[M]).$

### v3

@@ D.1.1 Experiment on L&O-NAE-SAT distribution
We consider five instances of L&O-NAE-SAT: (N,P)=(25,275),(30,270),(40,260),(50,250),(100,200)(N,P)=(25,275),(30,270),(40,260),(50,250),(100,200). For each distribution, we train a 19M MDM and measure the accuracy difference between vanilla inference and adaptive inference using top probability margin.
@@ D.1.2 Experiment on text data
Top probability margin sampler with temperature.
To modify our inference for text data modeling, which does not have a determined answer, we found that adding a certain level of temperature to the oracle is useful. This is because the top probability margin or the top probability often leads to greedy sampling, which harms the diversity (entropy) of the generated samples. Therefore, we consider a variant of the oracle as follows, incorporating a Gaussian noise term $\epsilon$.
@@ D.2 Experimental details on Sudoku and Zebra puzzles
For the training and inference, we use the codebase of (Ye et al., 2024) with keeping most of the hyperparameters default given in the codebase. For the Sudoku dataset, we use 66M GPT-2 model, and for the Zebra dataset, we use 1919M model. We set the learning rate to 0.001 with a batch size of 128 to train the model for 300 epochs. For the inference, we use 50 reverse sampling steps using the appropriate strategy. Additionally, we add Gumbel noise with a coefficient of 0.5 to the MDM inference oracle $\mathcal{F}$.
D.3 Experimental details on LLaDA-8B
Our evaluation covers two task categories: (i) infilling(HumanEval-Infill and ROCStories) and (ii) instruction–answering (Math). For instruction–answering tasks, we employ a semi-autoregressive sampling strategy, whereas for infilling tasks we retain the non-autoregressive approach. For infilling tasks, the output length is predetermined—matching the size of the masked span—whereas instruction–answering tasks require an explicit length specification. For the latter, we follow the sampling configuration of (Nie et al., 2025).
For HumanEval-Infill, we adopt the problem set introduced by (Bavarian et al., 2022). Each instance is grouped by the span of the masked code—the region the model must infill—into three categories: single-line, multi-line, and split. The task difficulty rises as the length of the masked span increases.
@@ Appendix E Omitted proofs
We build on Proposition 3.1 from (Zheng et al., 2024) to obtain the result of Proposition 2.1. We first re-state the result from (Zheng et al., 2024) for the case when the denoising network $p_{\theta}$ does not depend on the noise-scale $t$ explicitly. Let $x(n)$ be a sequence with $n$ tokens being masked from $x_{0}$, and $x^{i}(n)$ denotes the $i^{\textrm{th}}$ token value of the sequence $x(n)$. Let $\tilde{q}(x(n)|x_{0})$ be the probability distribution corresponding to randomly and uniformly masking $n$ tokens of $x_{0}$.
@@
For clean data $x_{0}$, let $\tilde{q}(x(n)\mid x_{0})$ be the discrete forward process that randomly and uniformly masks $n$ tokens of $x_{0}$. Suppose the noise schedules $\alpha_{t}$ satisfies $\alpha_{0}=0$ and $\alpha_{1}=1$. Then, the MDM training loss (1) can be reformulated as
$\mathcal{L}_{\theta}=-\sum_{n=1}^{L}\;\;\mathop{\mathbb{E}}_{x(n)\sim\tilde{q}(\cdot\mid x_{0})}\left[\frac{1}{n}\sum_{\ell:x^{\ell}(n)=0}\log p_{\theta}(x_{0}^{\ell}\mid x(n))\right].$ (6)
To obtain an alternative formulation of (6), we expand the expectation $x(n)\sim\tilde{q}(\cdot\mid x_{0})$. Since there are total $L$ positions of $x_{0}$, we have the probability assigned for each $x(n)$ equals $1/\binom{L}{n}$. Therefore, expanding the above equation with the expectation $x(n)$ and treating $x(n)$ as $x[M]$ for some set $M$ of size $n$, we obtain the result.
$\mathcal{L}_{\theta}=-\sum_{M\in[L],i\in M}\frac{1}{\binom{L}{|M|}}\cdot\frac{1}{|M|}\log p_{\theta}(x_{0}^{\ell}\mid x[M]).$
E.1 Equivalence between the MDM loss and any-order autoregressive loss
In this section, we will demonstrate the equivalence for MDM loss and any-order autoregressive loss. In particular, for all $x_{0}$, we show
$-\mathop{\mathbb{E}}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L})}\left[\sum_{j=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(j)}\Big{|}x_{0}[\pi\{j\},\ldots,\pi\{L-1\}]\right)\right]=-\sum_{M\subseteq[L],i\in M}\frac{1}{\binom{L}{|M|}}\frac{1}{|M|}\displaystyle\log p_{\theta}(x^{i}_{0}|x_{0}[M]).$
We now consider $\{\pi(j),\dots,\pi(L-1)\}=M\subseteq[L]$ and $\pi(j)=i$ and count the number of $\pi\in\mathbb{S}_{L}$ that induces a specific term $\log p_{\theta}(x_{0}^{i}|x_{0}[M])$. To induce the term, for a given $M\in[L]$ and $i\in M$, $\pi$ must satisfy
@@ E.1 Equivalence between the MDM loss and any-order autoregressive loss
The number of $\pi$ that satisfies above is $(L-|M|)!\times(|M|-1)!$. Using this and the number of total permutations is $L!$, we obtain the result.
$\mathop{\mathbb{E}}_{\pi\sim\mathrm{Unif}(\mathbb{S}_{L})}\left[\sum_{j=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(j)}\Big{|}x_{0}[\pi\{j\},\ldots,\pi\{L-1\}]\right)\right]$
$=$ $\frac{1}{L!}\sum_{\pi\in\mathrm{Unif}(\mathbb{S}_{L})}\sum_{j=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(j)}\Big{|}x_{0}[\pi\{j\},\ldots,\pi\{L-1\}]\right)$
$=$ $\frac{1}{L!}\sum_{M\in[L],i\in M}\big{[}\log p_{\theta}(x_{0}^{i}|x_{0}[M])\times(L-1-|M|)!\times(|M|-1)!\big{]}$
$=$ $\sum_{M\in[L],i\in M}\frac{1}{\binom{L}{|M|}}\frac{1}{|M|}\log p_{\theta}(x_{0}^{i}|x_{0}[M]).$